OpenStudy (anonymous):

Charlotte has two job offers as a car sales person. The first job will pay her $600 a month plus 2% on the price of her car sales. The second job will pay $1000 a month plus 1% on the price of her sales. How much must she sell each month to make the same income from the two jobs?

OpenStudy (anonymous):

600 + 0.02*N = 1000 + 0.01*N => 0.01N = 400 => N = 400/0.01 => N = $40,000